# Criteria for optimal entanglement-assisted long baseline imaging protocols

### v1

@@ Appendix D Discussion on different entanglement-assisted protocol
1 Single particle entangled states
@@
2 N-copy single particle entangled states
@@
In [15], a protocol is presented that uses only linear optics to estimate the phase $\theta$ of a point source (with $|g|=1$). This protocol achieves a Fisher information ratio for estimating $\mathbb{f}^{N}_{\theta}$ that scales as $1-\frac{1}{N+1}$ (defined as $\mathbb{F}{\theta}/\mathbb{H}{\theta}$), where $N$ is the number of copies of the shared single-particle entangled state (SPE). In this work, we extend their results by explicitly computing the Fisher information for estimating both $|g|$ and $\theta$ using their scheme for $N\in{1,\cdots,5}$.
$\mathbb{F}_{\theta}^{1}=\frac{1}{2}\frac{\sin^{2}\phi}{1-|g|^{2}\cos^{2}\phi}|g|^{2}\epsilon$ $\quad\mathbb{F}_{|g|}^{1}=\frac{1}{2}\frac{\cos^{2}\phi}{1-|g|^{2}\cos^{2}\phi}\epsilon$
$\mathbb{F}_{\theta}^{2}=\frac{3\sin^{2}\phi}{(1-|g|\cos\phi)(5+4|g|\cos\phi)}|g|^{2}\epsilon$ $\quad\mathbb{F}_{|g|}^{2}=\frac{3\cos^{2}\phi}{(1-|g|\cos\phi)(5+4|g|\cos\phi)}\epsilon$
$\mathbb{F}_{\theta}^{3}=\frac{3\sin^{2}\phi(9+7|g|\cos\phi)}{4(1-|g|^{2}\cos^{2}\phi)(10+6|g|\cos\phi)}|g|^{2}\epsilon$ $\quad\mathbb{F}_{|g|}^{3}=\frac{3\cos^{2}\phi(9+7|g|\cos\phi)}{4(1-|g|^{2}\cos^{2}\phi)(10+6|g|\cos\phi)}\epsilon$
$\mathbb{F}_{\theta}^{4}=\frac{10\sin^{2}\phi(16+9|g|\cos\phi)}{(1-|g|\cos\phi)(13+12|g|\cos\phi)(17+8|g|\cos\phi)}|g|^{2}\epsilon$ $\quad\mathbb{F}_{|g|}^{4}=\frac{10\cos^{2}\phi(16+9|g|\cos\phi)}{(1-|g|\cos\phi)(13+12|g|\cos\phi)(17+8|g|\cos\phi)}\epsilon$
$\mathbb{F}_{\theta}^{5}=\frac{5\sin^{2}\phi(79+106|g|\cos\phi+31|g|^{2}\cos^{2}\phi)}{(1-|g|^{2}\cos^{2}\phi)(26+10|g|\cos\phi)(20+16|g|\cos\phi)}|g|^{2}\epsilon$ $\quad\mathbb{F}_{|g|}^{5}=\frac{5\cos^{2}\phi(79+106|g|\cos\phi+31|g|^{2}\cos^{2}\phi)}{(1-|g|^{2}\cos^{2}\phi)(26+10|g|\cos\phi)(20+16|g|\cos\phi)}\epsilon$ (34)
@@
where $\phi=\delta-\theta$ with $\delta$ being the phase delay used in the measurement. In [15], they showed implicitly that $\phi=0$ is the optimal point to get the optimal Fisher information ratio $\mathbb{f}^{N}_{\theta}=1-\frac{1}{N+1}$ for point source with $|g|=1$, agree with Eq. 34 above, however, we could easily check that $\mathbb{f}^{N}_{\theta}=0$ at $\phi=0$ for a source with $|g|<1$ (which is almost always the case), in fact the Fisher information ratio scaling $1-\frac{1}{N+1}$ is an artifact that occurs because $|g|=1,\phi=0$ is a removable singular point. As some concrete examples, one can simply verify that:
$\mathbb{F}_{\theta}^{2}$ $=\frac{3\sin^{2}\phi}{(1-|g|\cos\phi)(5+4|g|\cos\phi)}|g|^{2}\epsilon\stackrel{{\scriptstyle\text{$|g|=1$, $\phi\rightarrow 0$}}}{{\Longrightarrow}}\frac{3\phi^{2}}{(\frac{1}{2}\phi^{2})(5+4)}\epsilon=\frac{2}{3}|g|^{2}\epsilon$ (35)
$\mathbb{F}_{\theta}^{2}$ $=\frac{3\sin^{2}\phi}{(1-|g|\cos\phi)(5+4|g|\cos\phi)}|g|^{2}\epsilon\stackrel{{\scriptstyle\text{$|g|<1$, $\phi\rightarrow 0$}}}{{\Longrightarrow}}0$ (36)
$\mathbb{F}_{\theta}^{2}$ $=\frac{3\sin^{2}\phi}{(1-|g|\cos\phi)(5+4|g|\cos\phi)}|g|^{2}\epsilon\stackrel{{\scriptstyle\text{$|g|\rightarrow 0$, $\phi\rightarrow\frac{\pi}{2}$}}}{{\Longrightarrow}}\frac{3}{5}|g|^{2}\epsilon$ (37)
@@
Unfortunately, we don’t have a close form answer on the achievable Fisher information ratio $\mathbb{f}^{N}_{\theta}$ and $\mathbb{f}^{N}_{|g|}$ using their scheme, because the computational complexity of their protocols (And actualy, any protocols that involves $N$-copy SPE states and linear optical circuit $\mathcal{U}$) is essentially related to the computational complexity of Boson sampling[53]. However, from our observable, we conject that:
@@
One could plug in $\phi=0$ for $\mathbb{f}_{|g|}^{N}$ and $\phi=\pi/2$ for $\mathbb{f}_{\theta}^{N}$ to Eq. 34 to check our conjecture.
3 KLM and modified KLM schemes
@@
$|\psi\rangle_{a}=\sum_{n=0}^{N}f_{n}|n_{A},(N-n)_{B}\rangle=\sum_{n=0}^{N}f_{n}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$ (38)
@@
$\mathbb{h}[|\psi\rangle^{\text{Itri}}_{a}]=\frac{16}{N^{2}}\sum_{n=1}^{N/2}\frac{i(i-1)}{i+i-1}=\frac{16}{N^{2}}\sum_{n=1}^{N/2}(\frac{i}{2}-\frac{1}{4}-\frac{1}{4}\frac{1}{2i-1})=1-\frac{4}{N^{2}}\sum_{n=1}^{N/2}\frac{1}{2i-1}\stackrel{{\scriptstyle\text{Large {\hbox{N}}}}}{{\approx}}1-2\frac{\log(N)}{N^{2}},$ (40)
@@
$\mathbb{h}[|\psi\rangle^{\text{Atri}}_{a}]=\frac{48}{N(N^{2}+2)}\sum_{n=1}^{N/2}\frac{i^{2}(i-1)^{2}}{i^{2}+(i-1)^{2}}=\frac{48}{N(N^{2}+2)}\sum_{n=1}^{N/2}(\frac{i^{2}}{2}-\frac{i}{2}-\frac{1}{4}+\frac{1}{4}\frac{1}{2i^{2}-2i+1})\stackrel{{\scriptstyle\text{Large {\hbox{N}}}}}{{\approx}}1-12\frac{1}{N^{2}},$ (41)
@@ 4 Optimal-KLM scheme
$|\psi\rangle^{\text{Sub}}_{a}=\sqrt{\frac{2}{N+2}}\sum\limits_{n=0}^{N}\sin(\frac{n+1}{N+2}\pi)|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B},$ (42)
@@
$\mathbb{h}[\rho^{AB}_{a}]$ $=\frac{4}{N+2}\sum_{n=1}^{N}\frac{\sin^{2}(\frac{n+1}{N+2}\pi)\sin^{2}(\frac{n}{N+2}\pi)}{\sin^{2}(\frac{n+1}{N+2}\pi)+\sin^{2}(\frac{n}{N+2}\pi)}=\frac{1}{N+2}\sum_{n=1}^{N}\frac{[\cos(\frac{\pi}{N+2})-\cos(\frac{2n+1}{N+2}\pi)]^{2}}{1-\cos(\frac{\pi}{N+2})\cos(\frac{2n+1}{N+2}\pi)}$
$=\frac{\cos(\frac{\pi}{N+2})}{N+2}\sum_{n=1}^{N}[\cos(\frac{\pi}{N+2})-\cos(\frac{2n+1}{N+2}\pi)]-\frac{\sin^{2}(\frac{\pi}{N+2})}{N+2}\sum_{n=1}^{N}\frac{\cos(\frac{2n+1}{N+2}\pi)[\cos(\frac{1}{N+2}\pi)-\cos(\frac{2n+1}{N+2}\pi)]}{1-\cos(\frac{\pi}{N+2})\cos(\frac{2n+1}{N+2}\pi)}$ (43)
@@
$\sum_{n=1}^{N}\cos(\frac{2n+1}{N+2}\pi)f(n)\leq\sum_{n=1}^{N}\cos(\frac{2n+1}{N+2}\pi)\cos(\frac{\pi}{N+2})\leq-\cos^{2}(\frac{\pi}{N+2})$
@@
5 Two-mode squeezed vacuum state
@@ A TMSQ state is written as
6 Two particle entangled state, and NOON state
@@
7 Correlated coherent states
@@
$E(y)=\textrm{Tr}_{a}[U_{sa}^{A}(y)\otimes U_{sa}^{B}(y)\sum_{n,m=0}^{\infty}(P^{A}_{n}\otimes P^{B}_{m})\rho_{s}^{AB}\otimes\rho_{a}^{AB}(P^{A}_{n}\otimes P^{B}_{m})U_{sa}^{A\dagger}(y)\otimes U_{sa}^{B\dagger}(y)]$ (53)
@@
Therefore, for a state with PR properties, i.e., $\mathbb{h}>0$, it is possible to find a gap between FI-ratio with all LOCC and QFI-ratio. The correlated coherent states will be a good example to that.
@@ Appendix E Implementation with linear optical quantum teleportation
$|\psi\rangle^{AB}=\frac{1}{\sqrt{2}}(|0\rangle_{A}|1\rangle_{B}+e^{i\theta}|1\rangle_{A}|0\rangle_{B})\sum_{n=0}^{N}f_{n}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$ (55)
@@
with $c_{0}=\sum_{\pi\in P^{0}_{n}}\prod_{i=1}^{n}\omega^{{d}_{i}(\vec{s})\pi(i)}\quad\quad c_{1}=\sum_{\pi\in P^{1}_{n}}\prod_{i=0}^{n-1}\omega^{{d}_{i}(\vec{s})\pi(i)}$ (61)
@@
$|\psi\rangle_{n}^{B}=\sum_{\vec{s}}\mathcal{P}_{\phi(\vec{s})}\prescript{A}{}{\langle\Psi(\vec{s})|}\mathcal{F}_{N+1}|\psi\rangle_{n}^{AB}=\frac{1}{\sqrt{|f_{n}|^{2}+|f_{n-1}|^{2}}}\left(f_{n}b^{\dagger}_{0}\prod_{j=n+1}^{N}b^{\dagger}_{j}+e^{i\theta}f_{n-1}\prod_{j=n}^{N}b^{\dagger}_{j}\right)|\text{{0}}\rangle^{B}$ (63)
@@
$\rho_{k}^{B}=\frac{1}{|f_{n}|^{2}+|f_{n-1}|^{2}}\begin{pmatrix}|f_{n}|^{2}&f_{n}f^{*}_{n-1}g\\ f^{*}_{n}f_{n-1}g^{*}&|f_{n-1}|^{2}\end{pmatrix},\quad\quad\quad\text{Written in }\{|0\rangle_{B}|1_{n}\rangle_{B},|1\rangle_{B}|0_{n}\rangle_{B}\}\text{ basis.}$. (65)
Followed by a tilted beam splitter measurement characterized by projection to |±$|\pm\rangle_{n}\propto f_{n-1}|0\rangle_{B}|1_{n}\rangle_{B}+f_{n}e^{i\delta}|1_{n}\rangle_{B}|0\rangle_{B}$ with a tunable phase delay $\delta$, we have
$P^{n}_{\pm}$ $=p_{n}\langle\pm|\rho_{n}^{B}|\pm\rangle_{n}=\frac{|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}(1\pm|g|\cos(\theta-\delta))\quad n\in[1,\cdots,N]$ (66)
where $p_{n}$ is given in Eq. 57. One can hence, evaluate the Fisher information for the corresponding scheme as:
$\mathbb{F}_{|g|}$ $=\sum_{n,\pm}\frac{1}{P^{n}_{\pm}}\left(\frac{\partial P^{n}_{\pm}}{\partial|g|}\right)^{2}=\sum_{n=1}^{N}\frac{2|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}\frac{\cos(\theta-\delta)^{2}}{1-|g|^{2}\cos(\theta-\delta)^{2}}\epsilon+O(\epsilon^{2})\xrightarrow{\theta=\delta,\epsilon\ll 1}\sum_{n=1}^{N}\frac{2|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}\frac{\epsilon}{1-|g|^{2}}$ (67)
$\mathbb{F}_{|\theta|}$ $=\sum_{n,\pm}\frac{1}{P^{n}_{\pm}}\left(\frac{\partial P^{n}_{\pm}}{\partial|g|}\right)^{2}=\sum_{n=1}^{N}\frac{2|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}\frac{\sin(\theta-\delta)^{2}|g|^{2}}{1-|g|^{2}\cos(\theta-\delta)^{2}}\epsilon+O(\epsilon^{2})\xrightarrow{\theta=\delta+\frac{\pi}{2},\epsilon\ll 1}\sum_{n=1}^{N}\frac{2|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}|g|^{2}\epsilon$ (68)

### v2

@@ Appendix D Discussion on different entanglement-assisted protocol
D.1 Single particle entangled states
@@
D.2 N-copy single particle entangled states
@@
In [15], a protocol is presented that uses only linear optics to estimate the phase $\theta$ of a point source (with $|g|=1$). This protocol achieves a Fisher information ratio for estimating $\mathbb{f}^{N}_{\theta}$ that scales as $1-\frac{1}{N+1}$, where $N$ is the number of copies of the shared single-particle entangled state (SPE). In this work, we extend their results by explicitly computing the Fisher information for estimating both $|g|$ and $\theta$ using their scheme for $N\in[1,\cdots,5]$.
$\mathbb{F}_{\theta}^{1}=\frac{1}{2}\frac{\sin^{2}\phi}{1-|g|^{2}\cos^{2}\phi}|g|^{2}\epsilon$
$\mathbb{F}_{|g|}^{1}=\frac{1}{2}\frac{\cos^{2}\phi}{1-|g|^{2}\cos^{2}\phi}\epsilon$
$\mathbb{F}_{\theta}^{2}=\frac{3\sin^{2}\phi}{(1-|g|\cos\phi)(5+4|g|\cos\phi)}|g|^{2}\epsilon$
$\mathbb{F}_{|g|}^{2}=\frac{3\cos^{2}\phi}{(1-|g|\cos\phi)(5+4|g|\cos\phi)}\epsilon$
$\mathbb{F}_{\theta}^{3}=\frac{3\sin^{2}\phi(9+7|g|\cos\phi)}{4(1-|g|^{2}\cos^{2}\phi)(10+6|g|\cos\phi)}|g|^{2}\epsilon$
$\mathbb{F}_{|g|}^{3}=\frac{3\cos^{2}\phi(9+7|g|\cos\phi)}{4(1-|g|^{2}\cos^{2}\phi)(10+6|g|\cos\phi)}\epsilon$
$\mathbb{F}_{\theta}^{4}=\frac{10\sin^{2}\phi(16+9|g|\cos\phi)}{(1-|g|\cos\phi)(13+12|g|\cos\phi)(17+8|g|\cos\phi)}|g|^{2}\epsilon$
$\mathbb{F}_{|g|}^{4}=\frac{10\cos^{2}\phi(16+9|g|\cos\phi)}{(1-|g|\cos\phi)(13+12|g|\cos\phi)(17+8|g|\cos\phi)}\epsilon$
$\mathbb{F}_{\theta}^{5}=\frac{5\sin^{2}\phi(79+106|g|\cos\phi+31|g|^{2}\cos^{2}\phi)}{(1-|g|^{2}\cos^{2}\phi)(26+10|g|\cos\phi)(20+16|g|\cos\phi)}|g|^{2}\epsilon$
$\mathbb{F}_{|g|}^{5}=\frac{5\cos^{2}\phi(79+106|g|\cos\phi+31|g|^{2}\cos^{2}\phi)}{(1-|g|^{2}\cos^{2}\phi)(26+10|g|\cos\phi)(20+16|g|\cos\phi)}\epsilon$
@@
where $\phi=\delta-\theta$ with $\delta$ being the phase delay used in the measurement. In [15], they showed implicitly that $\phi=0$ is the optimal point to get the optimal Fisher information ratio $\mathbb{f}^{N}_{\theta}=1-\frac{1}{N+1}$ for point source with $|g|=1$, agree with Eq. D.2 above, however, we could easily check that $\mathbb{f}^{N}_{\theta}=0$ at $\phi=0$ for a source with $|g|<1$ (which is almost always the case), in fact the Fisher information ratio scaling $1-\frac{1}{N+1}$ is an artifact that occurs because $|g|=1,\phi=0$ is a removable singular point. As some concrete examples, one can simply verify that:
$\mathbb{F}_{\theta}^{2}$ $=\frac{3\sin^{2}\phi}{(1-|g|\cos\phi)(5+4|g|\cos\phi)}|g|^{2}\epsilon$
[…]
$\mathbb{F}_{\theta}^{2}$ $=\frac{3\sin^{2}\phi}{(1-|g|\cos\phi)(5+4|g|\cos\phi)}|g|^{2}\epsilon$
[…]
$\mathbb{F}_{\theta}^{2}$ $=\frac{3\sin^{2}\phi}{(1-|g|\cos\phi)(5+4|g|\cos\phi)}|g|^{2}\epsilon$
[…]
@@
Unfortunately, we don’t have a close form answer on the achievable Fisher information ratio $\mathbb{f}^{N}_{\theta}$ and $\mathbb{f}^{N}_{|g|}$ using their scheme, because the computational complexity of their protocols (And actualy, any protocols that involves $N$-copy SPE states and linear optical circuit $\mathcal{U}$) is essentially related to the computational complexity of Boson sampling[49]. However, from our observation, we conject that:
@@
One could plug in $\phi=0$ for $\mathbb{f}_{|g|}^{N}$ and $\phi=\pi/2$ for $\mathbb{f}_{\theta}^{N}$ to Eq. D.2 to check our conjecture.
D.3 KLM and modified KLM schemes
@@
$|\psi\rangle_{a}$ $=\sum_{n=0}^{N}f_{n}|n_{A},(N-n)_{B}\rangle$
$=\sum_{n=0}^{N}f_{n}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$ (34)
@@
$\mathbb{h}[|\psi\rangle^{\text{Itri}}_{a}]=\frac{16}{N^{2}}\sum_{n=1}^{N/2}\frac{i(i-1)}{i+i-1}=\frac{16}{N^{2}}\sum_{n=1}^{N/2}(\frac{i}{2}-\frac{1}{4}-\frac{1}{4}\frac{1}{2i-1})$
$=1-\frac{4}{N^{2}}\sum_{n=1}^{N/2}\frac{1}{2i-1}\stackrel{{\scriptstyle\text{Large $N$}}}{{\approx}}1-2\frac{\log(N)}{N^{2}},$ (36)
@@
$\mathbb{h}[|\psi\rangle^{\text{Atri}}_{a}]=\frac{48}{N(N^{2}+2)}\sum_{n=1}^{N/2}\frac{i^{2}(i-1)^{2}}{i^{2}+(i-1)^{2}}$
$=\frac{48}{N(N^{2}+2)}\sum_{n=1}^{N/2}(\frac{i^{2}}{2}-\frac{i}{2}-\frac{1}{4}+\frac{1}{4}\frac{1}{2i^{2}-2i+1})$
$\stackrel{{\scriptstyle\text{Large $N$}}}{{\approx}}1-12\frac{1}{N^{2}},$ (37)
@@ D.4 Optimal-KLM scheme
$|\psi\rangle^{\text{Sub}}_{a}=\frac{\sum\limits_{n=0}^{N}\sin(\frac{n+1}{N+2}\pi)}{\sqrt{(N+2)/2}}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B},$ (38)
@@
$\mathbb{h}[\rho^{AB}_{a}]$ $=\frac{4}{N+2}\sum_{n=1}^{N}\frac{\sin^{2}(\frac{n+1}{N+2}\pi)\sin^{2}(\frac{n}{N+2}\pi)}{\sin^{2}(\frac{n+1}{N+2}\pi)+\sin^{2}(\frac{n}{N+2}\pi)}$
$=\frac{1}{N+2}\sum_{n=1}^{N}\frac{[\cos(\frac{\pi}{N+2})-\cos(\frac{2n+1}{N+2}\pi)]^{2}}{1-\cos(\frac{\pi}{N+2})\cos(\frac{2n+1}{N+2}\pi)}$
$=\frac{\cos(\frac{\pi}{N+2})}{N+2}\sum_{n=1}^{N}[\cos(\frac{\pi}{N+2})-\cos(\frac{2n+1}{N+2}\pi)]$
$-\frac{\sin^{2}(\frac{\pi}{N+2})}{N+2}\sum_{n=1}^{N}\frac{\cos(\frac{2n+1}{N+2}\pi)[\cos(\frac{1}{N+2}\pi)-\cos(\frac{2n+1}{N+2}\pi)]}{1-\cos(\frac{\pi}{N+2})\cos(\frac{2n+1}{N+2}\pi)}$
@@
$\sum_{n=1}^{N}\cos(\frac{2n+1}{N+2}\pi)f(n)$ $\leq\sum_{n=1}^{N}\cos(\frac{2n+1}{N+2}\pi)\cos(\frac{\pi}{N+2})$
$\leq-\cos^{2}(\frac{\pi}{N+2})$
@@
D.5 Two-mode squeezed vacuum state
@@ A TMSQ state is written as
D.6 Two particle entangled state, and NOON state
@@
D.7 Correlated coherent states
@@
$E(y)=\textrm{Tr}_{a}[U_{sa}^{A}(y)\otimes U_{sa}^{B}(y)\mathcal{E}(\rho^{AB})U_{sa}^{A\dagger}(y)\otimes U_{sa}^{B\dagger}(y)]$
$\mathcal{E}(\rho^{AB})=\sum_{n,m=0}^{\infty}(P^{A}_{n}\otimes P^{B}_{m})\rho_{s}^{AB}\otimes\rho_{a}^{AB}(P^{A}_{n}\otimes P^{B}_{m})$
@@
Therefore, for a state with PR properties, i.e., $\mathbb{h}>0$, it is possible to find a gap between FI-ratio with all LOCC and QFI-ratio. The correlated coherent states is an example to that.
@@ Appendix E Implementation with linear optical quantum teleportation
$|\psi\rangle^{AB}$ $=\frac{1}{\sqrt{2}}(|0\rangle_{A}|1\rangle_{B}+e^{i\theta}|1\rangle_{A}|0\rangle_{B})$
$\otimes\sum_{n=0}^{N}f_{n}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$ (49)
@@
with $c_{0}=\sum_{\pi\in P^{0}_{n}}\prod_{i=1}^{n}\omega^{{d}_{i}(\vec{s})\pi(i)}$ and $c_{1}=\sum_{\pi\in P^{1}_{n}}\prod_{i=0}^{n-1}\omega^{{d}_{i}(\vec{s})\pi(i)}$
@@
$|\psi\rangle_{n}^{B}$ $=\sum_{\vec{s}}\mathcal{P}_{\phi(\vec{s})}\prescript{A}{}{\langle\Psi(\vec{s})|}\mathcal{F}_{N+1}|\psi\rangle_{n}^{AB}$
$=\frac{\left(f_{n}b^{\dagger}_{0}\prod_{j=n+1}^{N}b^{\dagger}_{j}+e^{i\theta}f_{n-1}\prod_{j=n}^{N}b^{\dagger}_{j}\right)|\text{{0}}\rangle^{B}}{\sqrt{|f_{n}|^{2}+|f_{n-1}|^{2}}}$ (56)
@@
$\rho_{k}^{B}=\frac{1}{|f_{n}|^{2}+|f_{n-1}|^{2}}\begin{pmatrix}|f_{n}|^{2}&f_{n}f^{*}_{n-1}g\\ f^{*}_{n}f_{n-1}g^{*}&|f_{n-1}|^{2}\end{pmatrix},$ (58)
d which is written in $\{|0\rangle_{B}|1_{n}\rangle_{B},|1\rangle_{B}|0_{n}\rangle_{B}\}$ basis.. Followed by a tilted beam splitter measurement characterized by projection to |±$|\pm\rangle_{n}\propto f_{n-1}|0\rangle_{B}|1_{n}\rangle_{B}+f_{n}e^{i\delta}|1_{n}\rangle_{B}|0\rangle_{B}$ with a tunable phase delay $\delta$, we have
$P^{n}_{\pm}$ $=p_{n}\langle\pm|\rho_{n}^{B}|\pm\rangle_{n}=\frac{|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}(1\pm|g|\cos(\theta-\delta)),$ (59)
for $n\in[1,\cdots,N]$, where $p_{n}$ is given in Eq. 51. One can hence, evaluate the Fisher information for the corresponding scheme as:
$\mathbb{F}_{|g|}$ $=\sum_{n,\pm}\frac{1}{P^{n}_{\pm}}\left(\frac{\partial P^{n}_{\pm}}{\partial|g|}\right)^{2}$
$=\sum_{n=1}^{N}\frac{2|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}\frac{\cos(\theta-\delta)^{2}}{1-|g|^{2}\cos(\theta-\delta)^{2}}\epsilon+O(\epsilon^{2})$
$\xrightarrow{\theta=\delta,\epsilon\ll 1}\sum_{n=1}^{N}\frac{2|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}\frac{\epsilon}{1-|g|^{2}}$ (60)
$\mathbb{F}_{|\theta|}$ $=\sum_{n,\pm}\frac{1}{P^{n}_{\pm}}\left(\frac{\partial P^{n}_{\pm}}{\partial|g|}\right)^{2}$
$=\sum_{n=1}^{N}\frac{2|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}\frac{\sin(\theta-\delta)^{2}|g|^{2}}{1-|g|^{2}\cos(\theta-\delta)^{2}}\epsilon+O(\epsilon^{2})$
$\xrightarrow{\theta=\delta+\frac{\pi}{2},\epsilon\ll 1}\sum_{n=1}^{N}\frac{2|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}|g|^{2}\epsilon$ (61)